OpenStudy (anonymous):

Ok here is another one *ducks* dont shoot me! I cant help it that i suck at this stuff XD

OpenStudy (anonymous):

Ok it says to multiply then simplify if possible, let me type it out now

OpenStudy (anonymous):

\[\sqrt{10}(\sqrt{10}+z \sqrt{30})\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{10}(\sqrt{10}+z\sqrt{30})\] \[\Large \sqrt{10}\sqrt{10}+z\sqrt{10}\sqrt{30}\] \[\Large \sqrt{10*10}+z\sqrt{10*30}\] \[\Large \sqrt{100}+z\sqrt{300}\] \[\Large \sqrt{100}+z\sqrt{100*3}\] \[\Large \sqrt{100}+z\sqrt{100}*\sqrt{3}\] \[\Large 10+z*10*\sqrt{3}\] \[\Large 10+10z\sqrt{3}\] So \[\Large \sqrt{10}(\sqrt{10}+z\sqrt{30})=10+10z\sqrt{3}\]

OpenStudy (anonymous):

Thanks so much, your a life saver! :D How did you get the squared three though? that part lost me

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{300}\] \[\Large \sqrt{100*3}\] \[\Large \sqrt{100}*\sqrt{3}\] \[\Large 10\sqrt{3}\] So \[\Large \sqrt{300}=10\sqrt{3}\]

OpenStudy (anonymous):

So you just took the square of 300?

jimthompson5910 (jim_thompson5910):

I simplified the square root of 300

jimthompson5910 (jim_thompson5910):

taking the square root of 300 will give me a decimal number